Use the given data values​ (a sample of female arm circumferences in​ centimeters) to identify the corresponding z scores that are used for a normal quantile​ plot, then identify the coordinates of each point in the normal quantile plot. Construct the normal quantile​ plot, then determine whether the data appear to be from a population with a normal distribution.

38.8​, 32.2​, 34.1​, 40.1​, 45.0
Answer #1
i Sample quantiles=Sort(x) P=i/(n+1) Theoritical Quantiles=z=Norminv(P)
1 32.2 0.166667 -0.96742157
2 34.1 0.333333 -0.4307273
3 38.8 0.5 0
4 40.1 0.666667 0.430727299
5 45 0.833333 0.967421566

QQ plot Theoretical Quantiles 10 20 30 40 50 Sample Quantiles

The above QQ-plot shows a near to straight line by the dots. Hence, we can say that the data appears to be from a population with a normal distribution.